Simplify the trigonometric expression.
cos^3(x)+ sin^2(x)cos(x)

Answer by stanbon(48558)

Simplify the trigonometric expression.
cos^3(x)+ sin^2(x)cos(x)
-------------
Common factor: cos(x)
------------------
= (cos(x))(cos^2(x) + sin^2(x))
---
= cos(x)(1)
= cos(x)
=================
Cheers,
Stan H.